A tank that is a right rectangular prism is 1m x 2m x 3m. Suppose the tank is exactly half full of water and then a heavy metal sphere of radius 30 cm is put into the tank. How high is the water now if the height of the tank is 3m?
Answer by Alan3354(6097)

A tank that is a right rectangular prism is 1m x 2m x 3m. Suppose the tank is exactly half full of water and then a heavy metal sphere of radius 30 cm is put into the tank. How high is the water now if the height of the tank is 3m?
--------------
The volume of water in the tank is the area of the bottom x depth of the water.
= 100*200 sq cm * depth
The volume of the sphere
= 4*pi*r^3/3 cc
= 113097 cc
--------------
The tank is 1/2 full, then 113097 cc are added.
V1 = 100*200*150 cc = 3000000 cc
+ 113097 --> 3113097 cc
Depth = 3113097/20000
Depth = 155.65 cm